OpenStudy (anonymous):

please help me solve ln 3 − 2 ln 7 + ln 14

OpenStudy (anonymous):

PEMDAS

OpenStudy (anonymous):

Parenthesis Exponets Multiply Divide Addition Subtraction

OpenStudy (anonymous):

there are no parenthesis

OpenStudy (anonymous):

So skip that

OpenStudy (anonymous):

And btw can you say your question again?

OpenStudy (anonymous):

\[\ln 3-2\ln 7+\ln 14 \] I need to write the expression as a single logarithm with a coefficient of 1.